solve for x and y

x=12-y 5x+2y=12

Answer by checkley71(8403)

x=12-y now substitute (12-y) for x in the other eqaution & solve for y
5(12-y)+2y=12
60-5y+2y=12
-3y=12-60
-3y=-48
y=-48/-3
y=16 answer.
x=12-16
x=-4 answer.
proof
5*-4+2*12=12
-20+24=12
12=12